OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3.

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (anonymous):

ohhh okay

pooja195 (pooja195):

Can you tell the answer you got

OpenStudy (anonymous):

yes, thank you so much